3b.
this one has me simply dumbfounded. Please help if you can. Thanks.
How many solutions does the following system have:
4x + 8y = 16
3x + 6y = 9
Explain how to determine the number of solutions without solving the system. Then apply elimination, and interpret the resulting expression.

Answer by stanbon(75887)

How many solutions does the following system have:
4x + 8y = 16
3x + 6y = 9
Explain how to determine the number of solutions without solving the system. Then apply elimination, and interpret the resulting expression.
----------
Put each equation in slope-intercept form:
y = (-1/2)x + 2
y = (-1/2)x + (3/2)
-----------------------
They have the same slope and different y-intercepts so there are zero solutions.
---------------------------
By elimination:
4x + 8y = 16
3x + 6y = 9
---
Multiply thru the 2nd Eq. by (4/3) to get:
4x + 8y = 16
4x + 8y = 12
-----------------
Subtracting 2nd by 1st you get:
0 = 4
The system is inconsistent and has no solutions.
==============================
Cheers,
Stan H.